A loan of $8,400 was received at a rate of 9% compounded quarterly. Payments will be made at the end of each month for 5 years. Determine the size of the payments. Simple interest is paid for part of a period.
Answer #1

Annual Rate = 9%, Compounding Frequency : Quarterly

Applicable Quarterly Rate = 9/4 = 2.25 %

Applicable Monthly Rate = (1.0225)^(1/4) - 1 = 0.00558 or 0.558 %

Loan = $ 8400 and Tenure = 5 years or ( 5 x 12) = 60 months

Let the monthly repayments be $ k

Therefore, 8400 = k x (1/0.00558) x [1-{1/(1.00558)^(60)}]

8400 = k x 50.8723

k = 8400 / 50.8723 = $ 165.119 ~ $ 165.12